Suppose you roll a fair, six–faced die once. Find the probability of the event that the face 2 appears condition on the event that an even number has already occurred.
Answer #1

P(face 2 appears | an even number has occurred) = P(face 2 appears and an even number has occurred)/P(an even number has occurred) = (1/6)/(3/6) = 1/3 = 0.3333